OpenStudy (anonymous):

what is 3 2/5 + 2 4/5 in simplest form

hero (hero):

For your information: 3 2/5 = 3 + 2/5 2 4/5 = 2 + 4/5 Grouping like terms: (3 + 2) + 2/5 + 4/5 = 5 + 2/5 + 4/5 Now 2/5 = 1/5 + 1/5 so 5 + 1/5 + 1/5 + 4/5 = 5 + 1/5 + 5/5 5/5 = 1 so 5 + 1/5 + 5/5 = 5 + 1/5 + 1 If you understood everything you can finish the rest yourself.